OpenStudy (anonymous):

\[n \ge 2,\]\[\sum_{k=0}^{n-1}(x^{2n})^k=g_{n}(x)\sum_{k=0}^{n-1}x^k+n\]\[Find\ \ g_{n}(x)\ \ which\ is\ a\ polynomial \ with \ real \ coefficients.\]

OpenStudy (shubhamsrg):

well both are GPs (ones under summation sign) i hope this'll help !

OpenStudy (shubhamsrg):

<ding> ?

OpenStudy (anonymous):

i can't get the polynomial.

OpenStudy (shubhamsrg):

leme try.. we have ((x^2(n^2)) -1)/(x^2n -1) = gx (x^n -1)/(x-1) + n i'd guess just solve for gx..not pretty sure if it can be simplified further..

OpenStudy (anonymous):

but i can't get the polynomial form. i just get\[\frac{(1-x) (1-(x^{2n})^n+n (-1+x^{2n}))}{(-1+x^n) (-1+x^{2n})}\]

OpenStudy (shubhamsrg):

ohh i see..i'll tell you in some time if i can help or not :P

OpenStudy (experimentx):

what did you do to n??

OpenStudy (anonymous):

this?\[\frac{(1-x) (1-(x^{2n})^n+n (-1+x^{2n}))}{(-1+x^n) (-1+x^{2n})}\]

OpenStudy (shubhamsrg):

lol @experimentX ,,buddy dont worry,,he has done it correctly..

OpenStudy (experimentx):

lol ... prof, can you show the trick a bit??

OpenStudy (anonymous):

i think it is not correct

OpenStudy (anonymous):

I am still verifying it

OpenStudy (anonymous):

May be, stay tuned.

OpenStudy (anonymous):

Here it is \[ (x-1) \left(x^n+1\right) \sum _{k=1}^n (n-k) x^{2 (k-1) n} \]

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

how to get it?

OpenStudy (anonymous):

\[ f[x_]=\sum_{i=0}^n x^i \\ f(x^{2n})=\left((x-1) \left(x^n+1\right) \sum _{k=1}^n (n-k) x^{2 (k-1) n}\right) f(x) + n \]

OpenStudy (anonymous):

If you read my proof of the related problem I posted, you can realize it as \[ (x-1) h(x^n) \]

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

Oh ... i guess i got it too \[ x^{2 (k-2) n}\]